# Generating multipartite nonlocality to benchmark quantum computers

### v2

@@ III.2 Random sampling
So far, we have noted that there are multipartite Bell inequalities known that adapt to the architecture of the quantum computer and show a growing resistance to white noise. This allows to overcome problem (i) and (ii). We have seen however that also the number of terms grows exponentially, which makes it experimentally infeasible to measure all terms. In this section, we describe a scheme to overcome this problem and which allows to evaluate the Bell inequalities.
A general Bell operator $\mathcal{B}$ can be written as a sum of observables:
@@
We note that for the Bell inequalities in Sec. II the number of terms equals the quantum bound, i.e., $M=Q_{n}$. We propose to estimate the expectation value of the Bell operator by measuring the expectation value of $L$ randomly chosen terms. To analyse how many terms $L$ have to be sampled, we first assume that the expectation values can be inferred directly, i.e., we consider the limit of infinite measurements. With this simplification, the estimator reads
@@
We note that in the above expression $J_{1},\ldots,J_{L}$ are independent random variables with possible outcomes in the range $\{1,\ldots,M\}$. We assume that all outcomes are equally likely, i.e., $\mathbb{P}(J_{l}=j)=\frac{1}{M}$ for all $j\in\{1,\ldots,M\}$. In App. A.1, we show that the estimator is unbiased, i.e., $\mathbb{E}[\langle{\hat{\mathcal{B}}}\rangle_{\infty}]=\langle{\mathcal{B}}\rangle$.
To assess the significance of an observed violation, we consider the $p$-value. We upper bound the $p$-value with the help of Eq. (11). The probability on the right hand side of Eq. (11) can be bounded with the help of concentration inequalities. Here, we consider Hoeffding’s inequality, which yields
@@
The details on Hoeffding’s inequality are presented in App. B.1. This result can be used to derive a lower bound on the necessary $L$. To reach a confidence level of $\gamma=1-p$,
$L\geq\left\lceil-\frac{2M^{2}}{t^{2}}\log(1-\gamma)\right\rceil$ (15)
@@ III.3 Number of measurement repetitions
We now take into account that the expectation values cannot be inferred directly. Rather, the measurement of each chosen term $B_{j}$ has to be repeated multiple times. In this section, we are interested in how many repetitions are necessary. For this purpose, we adjust the estimator in Eq. (13) to include the measurement repetitions. We assume that every term is measured $K$-times. This yields the estimator
@@
In the above estimator, $b_{j}^{(k)}$ denotes the measurement outcome of the term $B_{j}$ in the $k$-th repetition. Thus, we have $b_{j}^{(k)}\in\{\pm 1\}$. As in Eq. (13), $J_{1},\ldots,J_{L}$ denote independent random variables that uniformly distributed take values from $\{1,\ldots,M\}$. We show in App. A.2 that also the estimator in Eq. (16) is unbiased, i.e., $\mathbb{E}[\langle{\hat{\mathcal{B}}}\rangle]=\langle{\mathcal{B}}\rangle$. Finally, we can again use Eq. (11) and Hoeffding’s inequality to bound the $p$-value:
@@
$K\geq\frac{1}{L}\left\lceil-\frac{2M^{2}}{t^{2}}\log(1-\gamma)\right\rceil.$ (18)
Since $L\geq\lceil-\frac{2M^{2}}{t^{2}}\log(1-\gamma)\rceil$, we observe that $K\geq 1$. As a result, we can conclude that it is sufficient to measure each term only once.
Figure 3: Necessary number of random observables $L$, which are measured $K=1$ times each, such that an observed violation of at least ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ for $\alpha=0.6$ has a confidence of $\gamma=5\sigma$. $L$ is plotted as a function of the number of qubits $n$.
As an example, we consider the case that a violation of at least ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ for $\alpha=0.6$ is observed. This value is motivated by Ref. [36], where they prepared a $51$ qubit LC state with fidelity of approximately $0.6$. The necessary number of mesurement settings $L$ is shown in Fig. 3. We note that for small $n$, $L$ exceeds the number of contexts of the Bell operator. This, however, is not a contradiction as we do not exclude that a term is sampled multiple times. Exemplary we can make the observation:
@@ IV Analysis of the Bell inequalities for the GHZ and LC state
In this section, we are going to apply the method to the Bell inequalities for graph states. Especially, we will look at the Bell inequalities for the GHZ and the LC state. As described in the motivation, these Bell inequalities are promising to detect a large violation. The discussion in the previous section has shown that this is necessary to verify the violation from few measurement settings with high significance. We start with the question how the LC and GHZ state can be prepared on a quantum computer with a given two-qubit connectivity. Afterwards, we consider the effect of noise on the preparation with a simple depolarisation noise model, which gives insight into the sample complexity of the method.
@@ IV.1 Connectivities of current quantum computers
We start by having a look at the connectivity graphs of different quantum computers and the Bell inequalities that can be evaluated on the different architectures. In Fig. 5, we show the connectivity graphs of few current quantum computers. The first connectivity in Fig. 5a is the star graph of five qubits, i.e., one central qubit connected to four other qubits. This layout is used, e.g., in the Starmon-5 quantum processor [44]. Fig. 5b shows the connectivity graph that is used by IBM’s Falcon processor [45]. The ion trap quantum computer in Ref. [46] has 20-qubits that can all be coupled. The corresponding connectivity graph is shown in Fig. 5c. We also include the connectivity graphs of Google’s Sycamore processor in Fig. 5d [13] that has $53$ qubits and IBM’s Eagle processor [45] that has $127$ qubits.
Figure 4: Preparation scheme for the GHZ state on the $7$-qubit interaction topology that is for example used by IBM’s Falcon processor. The dotted lines indicate the physical CZ gates, whereas the graph state is drawn in red.
Figure 5: Connectivities of different quantum computers. The connectivity graph in (a) is used for example by the the Starmon-5 quantum processor [44], whereas IBM’s Falcon processor [45] is based on layout (b). (c) shows the connectivity of the ion trap quantum computer in Ref. [46]. Finally, in (d) we depict the connectivity graph of Google’s Sycamore processor [13] and in (e) IBM’s Eagle processor ibm_brisbane [45].
As the optimal Bell inequality that is associated to the connectivity graph is in general hard to determine, we will focus on the Bell inequalities for the GHZ and the LC state. In the following, we are interested in the largest GHZ and LC state that can be prepared on the different layouts in Fig. 5. Ideally, the preparation only requires the set of basis gates that are directly implemented by the quantum computer.
@@
The connectivity graph of quantum computers is usually connected. Thus, by the following steps a GHZ state of all $n$ qubits can be prepared. The steps are illustrated in Fig. 4 for the architecture in Fig. 5b.
  1. 1.
    Prepare a star graph with center at the qubit with the largest connectivity [Fig. 4b].
  2. 2.
    By performing two local complementations, the center of the star graph can be shifted to any node of the graph. Thus, the center can be moved to a vertex with still uncoupled neighbours [Fig. 4c-d].
  3. 3.
    By applying a CZ gate between the center node and the uncoupled neighbours the adjacent qubits can be added to the GHZ state [Fig. 4e].
  4. 4.
    Step 2+3 can be repeated until all qubits are coupled.
@@
For the LC state in contrast, we make the following observation.
@@
Assume that the connectivity graph of a quantum computer is connected. Then, a linear cluster state containing all $n$ qubits can be prepared with a circuit depth of $\mathcal{O}(n)$. In practice, however, it is often beneficial to prepare the LC state that is associated to the longest simple path in the connectivity graph [36]. The corresponding circuit has a constant depth of three independent of the number of qubits.
@@
We show in App. C that it is always possible to prepare a LC state that contains all qubits in linear circuit depth. The longest simple path, in contrast, can be generated by first preparing all qubits in the $\ket{+}$ state, i.e., by applying Hadamard gates. Afterwards, every second CZ gate can be performed in parallel. The circuit depth is thus three independent of the length of the simple path. The circuit is shown for the $6$-qubit LC state in Fig. 6a. ∎
As we only know the Bell inequality for the LC state with a number of qubits that is a multiple of three, we search for the longest path of length divisible by three. The longest paths that fulfill this restriction are drawn pink in Fig. 5. The connectivity graphs in Fig. 5a and Fig. 5b allow to prepare a $3$-qubit LC state. The largest simple path on the $20$-qubit full-connectivity graph in Fig. 5c with length divisible by three has length $18$. The quantum computer in Fig. 5c thus allows to check the Bell inequality for the $18$-qubit LC state. To find the longest path is an NP-complete problem [49]. We can thus not verify if the marked paths for the layouts in Fig. 5d and Fig. 5e are indeed the longest paths. In Fig. 5d, we have identified a $48$-qubit path as longest simple path. The layout in Fig. 5e, in turn, allows the preparation of the $108$-qubit LC state.
@@ IV.2 Noise
In the following section, we will discuss the effect of noise. The goal is to roughly estimate the violation that can be realistically observed on quantum computers and how the violation scales with the number of qubits $n$. Commonly, the errors of quantum computers are specified in terms of the error rates for single-qubit gates, two-qubit gates and readout. The average error rates for IBM’s Eagle processor ibm_brisbane and Google’s Sycamore processor are shown in Table 1. The error rate for the single-qubit gates is typically about an order smaller than the other error rates. The readout error on the other side can be mitigated by classical post-processing [50, 51]. To consider the effect of noise on the Bell violation, we will consider a simple depolarisation noise model.
@@
Table 1: Average error rates of IBM’s Eagle processor ibm_brisbane calibrated on 27.02.2024 18:32:14 [45] and Google’s Sycamore processor [13]. Google gives the error rates for the cases that the gates are performed isolated or simultaneously on all qubits. The error rates are averaged over all gates or the readout error of all qubits.
In the depolarisation noise model, we assume that an error results on average in a maximally mixed state, i.e., the circuit for the graph state |$\ket{G}$ prepares the mixture
@@
In Fig. 6, we show the circuits to prepare the LC and GHZ state in the case of $n=6$ qubits. For the LC state the CZ gates can be performed simultaneously and thus the circuit exhibits a constant depth of three independent of the number of qubits. The number of single-qubit gates is $N_{1}=n+2$, which includes two identity operations for the idling qubits. There are in total $n-1$ CZ gates, such that $N_{2}=n-1$. For the GHZ state, however, the preparation with CZ gates requires the gates to be applied consecutively. The circuit depth thus grows with the number of qubits. In each step, $n-2$ of the qubits are idling. Therefore, there are in total $(n-1)(n-2)$ identity operations, such that the preparation requires in total $N_{1}=n+(n-1)(n-2)$ single qubit gates. The number of CZ gates is equal, i.e., $N_{2}=n-1$.
@@
Table 2: Results for the depolarisation noise model. For the noise data of Google’s Sycamore [13] and IBM’s Eagle processor [45] we show the violations in terms of the quantum bound $Q$. The last column shows the number of random terms $L$ that have to be sampled to verify the violation with a confidence of $5\sigma$.
The results in Table 2 show that the simple noise model predicts a violation of approximately $0.1\times Q$ for Google’s Sycamore processor and around $0.02\times Q$ for ibm_brisbane. We note however that we considered more qubits on the IBM machine. Moreover, the violation of the GHZ state is smaller compared to the LC state. This indicates that the GHZ state is more affected by noise. A violation can accordingly be verified with $L\sim 2000$ measurement settings on Google’s Sycamore processor. On the IBM Eagle processor, in turn, $L\sim 40000$ settings are needed to verify nonlocality in the LC state and $L\sim 100000$ for the GHZ state. Real quantum computers, moreover, do not implement all gates natively. IBM’s Eagle processor for example does not support the CZ gate. Rather, it has to be composed by the available gates. The circuit in practice thus contains more gates and possibly exhibits a larger depth. For these reasons the noise model overestimates the violation. However, the noise model is still useful to assess the scaling of the sample complexity.
@@ IV.3 Sample complexity
Figure 7: Necessary number of random observables $L$, which are measured $K=1$ times each, such that an observed violation of at least ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ with $\alpha_{\text{GHZ}}=\exp(-an^{2})$ or $\alpha_{\text{LC}}=\exp(-bn)$ has a confidence of $\gamma=5\sigma$. $L$ is plotted as a function of the number of qubits $n$.
In the previous section, we have seen that noise causes the Bell violation to decline exponentially with the number of qubits, i.e., in leading order we have
@@
with $a,b>0$. As we have seen in Eq. (4), a violation of the Bell inequality can still be observed for large $n$ if $\alpha$ vanishes slower than the fraction $D_{n}^{-1}$, i.e., $\alpha>D_{n}^{-1}\overset{n\rightarrow\infty}{\rightarrow}0$. We note that for the LC state this is the case for $b<\log(2)/3$. In this parameter regime, the relative violation ^$\langle{\hat{\mathcal{B}}}\rangle/C$ of the Bell inequality for the LC state is still increasing with $n$. For the GHZ state, however, the scaling in Eq. (21) shows that the effect of noise increases faster than the quantum bound. A violation of the Bell inequality for the GHZ state can thus only be observed for $n<\ln(2)/(4a)+\sqrt{(\ln(2)/(4a))^{2}-\ln(2)/a}$. In case a violation is observed, it is $t=\alpha Q_{n}-C_{n}$. The number of random terms that is necessary to ensure a confidence level $\gamma$ is given by Eq. (15) and takes the form
$L\geq\left\lceil-\frac{2}{(\alpha-D_{n}^{-1})^{2}}\log(1-\gamma)\right\rceil.$ (22)
We note that, for the Bell inequalities in Sec. II, the number of terms equals the quantum bound, i.e., $M=Q_{n}$. In case $\alpha$ also decreases exponentially, the number of random settings $L$ increases exponentially. This is shown in Fig. 7.
For both the GHZ state as well as the LC state, we have plotted $L$ as a function of the number of qubits $n$. The scaling for the GHZ state is due to the number of single-qubit gates. We therefore choose the parameter $a$ of the same order as the error in the single-qubit gates, i.e., $a=0.001,0.0005$ and $0.0001$. For the LC state in contrast all errors contribute to the leading term and we choose $b=0.1,0.01$ and $0.001$. Fig. 7 shows that for $a\leq 0.0001$ or $b\leq 0.01$ the number of required measurements are feasible on current quantum computers.
@@
In addition, Eq. (22) shows that with increasing $L$ a decreasing violation ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ with $\alpha\sim\mathcal{O}(L^{-1/2})$ can be verified.
Finally, if the classical bound $C_{n}$ is negligible compared to the observed violation, it does not have an effect on $L$. We thus show the contours of the confidence levels $\gamma=1\sigma,2\sigma$ and $5\sigma$ in this limit for a given observation $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ with $L$ random terms in Fig. 8.

### v3

@@ III.2 Random sampling
So far, we have noted that there are known multipartite Bell inequalities that adapt to the architecture of the quantum computer and show a growing resistance to white noise. This allows one to overcome problems (i) and (ii). We have also seen, however, that the number of terms grows exponentially, which makes it experimentally infeasible to measure all terms. In this section, we describe a scheme to overcome this problem and which allows one to evaluate the Bell inequalities.
A general Bell operator $\mathcal{B}$ can be written as a sum of observables,
@@
We note that for the Bell inequalities in Sec. II, the number of terms equals the quantum bound, i.e., $M=Q_{n}$. We propose to estimate the expectation value of the Bell operator by measuring the expectation value of $L$ randomly chosen terms. To analyze how many terms $L$ have to be sampled, we first assume that the expectation values can be inferred directly, i.e., we consider the limit of infinite measurements. With this simplification, the estimator reads
@@
We note that in the above expression, $J_{1},\ldots,J_{L}$ are independent random variables with possible outcomes in the range $\{1,\ldots,M\}$. We assume that all outcomes are equally likely, i.e., $\mathbb{P}(J_{l}=j)=\frac{1}{M}$ for all $j\in\{1,\ldots,M\}$. In Appendix A.1, we show that the estimator is unbiased, i.e., $\mathbb{E}[\langle{\hat{\mathcal{B}}}\rangle_{\infty}]=\langle{\mathcal{B}}\rangle$.
To assess the significance of an observed violation, we consider the $p$ value. We upper bound the $p$ value with the help of Eq. (11). The probability on the right-hand side of Eq. (11) can be bounded with the help of concentration inequalities. Here, we consider Hoeffding’s inequality, which yields
@@
The details of Hoeffding’s inequality are presented in Appendix B.1. This result can be used to derive a lower bound on the necessary $L$. To reach a confidence level of $\gamma=1-p$,
$L\geq\left\lceil-\frac{2M^{2}}{t^{2}}\ln(1-\gamma)\right\rceil$ (15)
@@ III.3 Number of measurement repetitions
We now take into account that the expectation values cannot be inferred directly. Rather, the measurement of each chosen term $B_{j}$ has to be repeated multiple times. In this section, we are interested in how many repetitions are necessary. For this purpose, we adjust the estimator in Eq. (13) to include the measurement repetitions. We assume that every term is measured $K$ times. This yields the estimator
@@
In the above estimator, $b_{j}^{(k)}$ denotes the measurement outcome of the term $B_{j}$ in the $k$th repetition. Thus, we have $b_{j}^{(k)}\in\{\pm 1\}$. As in Eq. (13), $J_{1},\ldots,J_{L}$ denote independent random variables that when uniformly distributed, take values in the range $\{1,\ldots,M\}$. We also show in Appendix A.2 that the estimator in Eq. (16) is unbiased, i.e., $\mathbb{E}[\langle{\hat{\mathcal{B}}}\rangle]=\langle{\mathcal{B}}\rangle$. Finally, we can again use Eq. (11) and Hoeffding’s inequality to bound the $p$ value,
@@
$K\geq\frac{1}{L}\left\lceil-\frac{2M^{2}}{t^{2}}\ln(1-\gamma)\right\rceil.$ (18)
Since $L\geq\lceil-\frac{2M^{2}}{t^{2}}\ln(1-\gamma)\rceil$, we observe that $K\geq 1$. As a result, we can conclude that it is sufficient to measure each term only once.
Figure 3: Necessary number of random observables, $L$, which are measured $K=1$ times each, such that an observed violation of at least ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ for $\alpha=0.6$ has a confidence of $\gamma=5\sigma$. $L$ is plotted as a function of the number of qubits, $n$.
As an example, we consider the case that a violation of at least ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ for $\alpha=0.6$ is observed. This value is motivated by Ref. [36], where they prepared a $51$-qubit LC state with fidelity of approximately $0.6$. The necessary number of measurement settings, $L$, is shown in Fig. 3. We note that for small $n$, $L$ exceeds the number of contexts of the Bell operator. This, however, is not a contradiction as we do not exclude that a term is sampled multiple times. As an example, we can make the following observation:
@@ IV Analysis of the Bell inequalities for the GHZ and LC state
In this section, we are going to apply the method to the Bell inequalities for graph states. In particular, we will look at the Bell inequalities for the GHZ and the LC states. As described in the introduction, these Bell inequalities are promising to detect a large violation. The discussion in the previous section has shown that this is necessary to verify the violation from a few measurement settings with high significance. We start with the question of how the LC and GHZ states can be prepared on a quantum computer with a given two-qubit connectivity. Afterwards, we consider the effect of noise on the preparation with a simple depolarization noise model, which gives insight into the sample complexity of the method.
@@ IV.1 Connectivities of current quantum computers
We start by having a look at the connectivity graphs of different quantum computers and the Bell inequalities that can be evaluated on the different architectures. In Fig. 5, we show the connectivity graphs of a few current quantum computers. The first connectivity in Fig. 5(a) is the star graph of five qubits, i.e., one central qubit connected to four other qubits. This layout is used, e.g., in the Starmon-5 quantum processor [44]. Figure 5(b) shows the connectivity graph that is used by IBM’s Falcon processor [45]. The ion trap quantum computer in Ref. [46] has 20-qubits that can all be coupled. The corresponding connectivity graph is shown in Fig. 5(c). We also include the connectivity graphs of Google’s Sycamore processor in Fig. 5(d) [13] that has $53$ qubits and IBM’s Eagle processor [45] that has $127$ qubits in Fig. 5(e).
Figure 4: Preparation scheme for the GHZ state on the $7$-qubit interaction topology that is, for example, used by IBM’s Falcon processor. The dotted lines indicate the physical CZ gates, whereas the graph state is drawn in red.
Figure 5: Connectivities of different quantum computers. The connectivity graph in (a) is used, for example, by the the Starmon-5 quantum processor [44], whereas IBM’s Falcon processor [45] is based on layout (b). (c) The connectivity of the ion trap quantum computer in Ref. [46]. (d) The connectivity graph of Google’s Sycamore processor [13]; (e) IBM’s Eagle processor ibm_brisbane [45].
As the optimal Bell inequality that is associated to the connectivity graph is, in general, hard to determine, we will focus on the Bell inequalities for the GHZ and the LC states. In the following, we are interested in the largest GHZ and LC states that can be prepared on the different layouts in Fig. 5. Ideally, the preparation only requires the set of basis gates that are directly implemented by the quantum computer.
@@
The connectivity graph of quantum computers is usually connected. Thus, by the following steps, a GHZ state of all $n$ qubits can be prepared. The steps are illustrated in Fig. 4 for the architecture in Fig. 5(b).
  1. (1)
    Prepare a star graph with center at the qubit with the largest connectivity [Fig. 4(b)].
  2. (2)
    By performing two local complementations, the center of the star graph can be shifted to any node of the graph. Thus, the center can be moved to a vertex with still uncoupled neighbors [Figs. 4(c) and 4(d)].
  3. (3)
    By applying a CZ gate between the center node and the uncoupled neighbors the adjacent qubits can be added to the GHZ state [Fig. 4(e)].
  4. (4)
    Step (2) and (3) can be repeated until all qubits are coupled.
@@
For the LC state, in contrast, we make the following observation.
@@
Assume that the connectivity graph of a quantum computer is connected. Then, a linear cluster state containing all $n$ qubits can be prepared with a circuit depth of $\mathcal{O}(n)$. In practice, however, it is often beneficial to prepare the LC state that is associated to the longest simple path in the connectivity graph [36]. The corresponding circuit has a constant depth of three, independent of the number of qubits.
@@
We show in Appendix C that it is always possible to prepare a LC state that contains all qubits in linear circuit depth. The longest simple path, in contrast, can be generated by first preparing all qubits in the $\ket{+}$ state, i.e., by applying Hadamard gates. Afterwards, every second CZ gate can be performed in parallel. The circuit depth is thus three, independent of the length of the simple path. The circuit is shown for the $6$-qubit LC state in Fig. 6(a). ∎
As we only know the Bell inequality for the LC state with a number of qubits that is a multiple of three, we search for the longest path of length divisible by three. The longest paths that fulfill this restriction are drawn in pink in Fig. 5. The connectivity graphs in Figs. 5(a) and 5(b) allow one to prepare a $3$-qubit LC state. The largest simple path on the $20$-qubit full-connectivity graph in Fig. 5(c) with length divisible by three has length $18$. The quantum computer in Fig. 5(c) thus allows one to check the Bell inequality for the $18$-qubit LC state. To find the longest path is an NP-complete problem [49]. We can thus not verify if the marked paths for the layouts in Figs. 5(d) and 5(e) are indeed the longest paths. In Fig. 5(d), we have identified a $48$-qubit path as the longest simple path. The layout in Fig. 5(e), in turn, allows the preparation of the $108$-qubit LC state.
@@ IV.2 Noise
In the following section, we will discuss the effect of noise. The goal is to roughly estimate the violation that can be realistically observed on quantum computers and how the violation scales with the number of qubits, $n$. Commonly, the errors of quantum computers are specified in terms of the error rates for single-qubit gates, two-qubit gates, and readout. The average error rates for IBM’s Eagle processor ibm_brisbane and Google’s Sycamore processor are shown in Table 1. The error rate for the single-qubit gates is typically about an order smaller than the other error rates. The readout error on the other side can be mitigated by classical postprocessing [50, 51]. To consider the effect of noise on the Bell violation, we will consider a simple depolarization noise model.
@@
Table 1: Average error rates of IBM’s Eagle processor ibm_brisbane [45] and Google’s Sycamore processor [13]. Google gives the error rates for the cases that the gates are performed isolated or simultaneously on all qubits. The error rates are averaged over all gates or the readout error of all qubits.
In the depolarization noise model, we assume that an error results, on average, in a maximally mixed state, i.e., the circuit for the graph state |$\ket{G}$ prepares the mixture
@@
In Fig. 6, we show the circuits to prepare the LC and GHZ states in the case of $n=6$ qubits. For the LC state, the CZ gates can be performed simultaneously and thus the circuit exhibits a constant depth of three, independent of the number of qubits. The number of single-qubit gates is $N_{1}=n+2$, which includes two identity operations for the idling qubits. There are, in total, $n-1$ CZ gates, such that $N_{2}=n-1$. For the GHZ state, however, the preparation with CZ gates requires the gates to be consecutively applied. The circuit depth thus grows with the number of qubits. In each step, $n-2$ of the qubits are idling. Therefore, there are, in total, $(n-1)(n-2)$ identity operations, such that the preparation requires, in total, $N_{1}=n+(n-1)(n-2)$ single qubit gates. The number of CZ gates is equal, i.e., $N_{2}=n-1$.
@@
Table 2: Results for the depolarization noise model. For the noise data of Google’s Sycamore [13] and IBM’s Eagle processor [45], we show the violations in terms of the quantum bound $Q$. The last column shows the number of random terms, $L$, that have to be sampled to verify the violation with a confidence of $5\sigma$.
The results in Table 2 show that the simple noise model predicts a violation of approximately $0.1\times Q$ for Google’s Sycamore processor and around $0.02\times Q$ for ibm_brisbane. We note, however, that we considered more qubits on the IBM machine. Moreover, the violation of the GHZ state is smaller compared to the LC state. This indicates that the GHZ state is more affected by noise. A violation can accordingly be verified with $L\sim 2000$ measurement settings on Google’s Sycamore processor. On the IBM Eagle processor, in turn, $L\sim 40000$ settings are needed to verify nonlocality in the LC state and $L\sim 100000$ for the GHZ state. Real quantum computers, moreover, do not implement all gates natively. IBM’s Eagle processor, for example, does not support the CZ gate. Rather, it has to be composed of the available gates. The circuit in practice thus contains more gates and possibly exhibits a larger depth. For these reasons, the noise model overestimates the violation. However, the noise model is still useful to assess the scaling of the sample complexity.
@@ IV.3 Sample complexity
Figure 7: Necessary number of random observables $L$, which are measured $K=1$ times each, such that an observed violation of at least ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ with $\alpha_{\text{GHZ}}=\exp(-an^{2})$ or $\alpha_{\text{LC}}=\exp(-bn)$ has a confidence of $\gamma=5\sigma$. $L$ is plotted as a function of the number of qubits, $n$.
In the previous section, we have seen that noise causes the Bell violation to decline exponentially with the number of qubits, i.e., in leading order, we have
@@
with $a,b>0$. As we have seen in Eq. (4), a violation of the Bell inequality can still be observed for large $n$ if $\alpha$ vanishes slower than the fraction $D_{n}^{-1}$, i.e., $\alpha>D_{n}^{-1}\overset{n\rightarrow\infty}{\rightarrow}0$. We note that for the LC state, this is the case for $b<\ln(2)/3$. In this parameter regime, the relative violation ^$\langle{\hat{\mathcal{B}}}\rangle/C$ of the Bell inequality for the LC state is still increasing with $n$. For the GHZ state, however, the scaling in Eq. (21) shows that the effect of noise increases faster than the quantum bound. A violation of the Bell inequality for the GHZ state can thus only be observed for $n<\ln(2)/(4a)+\sqrt{(\ln(2)/(4a))^{2}-\ln(2)/a}$. In case a violation is observed, it is $t=\alpha Q_{n}-C_{n}$. The number of random terms that is necessary to ensure a confidence level $\gamma$ is given by Eq. (15) and takes the form
$L\geq\left\lceil-\frac{2}{(\alpha-D_{n}^{-1})^{2}}\ln(1-\gamma)\right\rceil.$ (22)
We note that for the Bell inequalities in Sec. II, the number of terms equals the quantum bound, i.e., $M=Q_{n}$. In case $\alpha$ also decreases exponentially, the number of random settings, $L$, increases exponentially. This is shown in Fig. 7.
For both the GHZ state as well as the LC state, we have plotted $L$ as a function of the number of qubits, $n$. The scaling for the GHZ state is due to the number of single-qubit gates. We therefore choose the parameter $a$ of the same order as the error in the single-qubit gates, i.e., $a=0.001,0.0005$, and $0.0001$. For the LC state, in contrast, all errors contribute to the leading term and we choose $b=0.1,0.01$, and $0.001$. Figure 7 shows that for $a\leq 0.0001$ or $b\leq 0.01$, the number of required measurements is feasible on current quantum computers.
@@
In addition, Eq. (22) shows that with increasing $L$, a decreasing violation ^$\langle{\hat{\mathcal{B}^{\text{}}}}\rangle=\alpha Q_{n}$ with $\alpha\sim\mathcal{O}(L^{-1/2})$ can be verified.
Finally, if the classical bound $C_{n}$ is negligible compared to the observed violation, it does not have an effect on $L$. We thus show the contours of the confidence levels $\gamma=1\sigma,2\sigma$, and $5\sigma$ in this limit for a given observation $\langle{\hat{\mathcal{B}^{\text{}}}}\rangle$ with $L$ random terms in Fig. 8.